# Differential privacy guarantees of Markov chain Monte Carlo algorithms

### v1

@@ 5 Privacy of Langevin-based algorithms in the non-convex setting
In this section we obtain (Rényi)-DP for both the trajectory and the final value of two Langevin-based MCMC algorithms: ULA and noisy SGD. All bounds in this section are derived from Propositions 4.5 and 4.6.
@@ 5.1 Sampling from Bayesian posteriors with ULA
with initial condition $x^{\mathcal{D}}_{0}=x_{0}\in\mathbb{R}^{d},$ and where $(z_{n})_{n\geqslant 1}$ is a sequence of i.i.d. standard Gaussians on $\mathbb{R}^{d}$, and $\gamma>0$ is the step size. ULA arises as the Euler discretisation of the overdamped Langevin diffusion and has been shown to be successful for sampling from Bayesian posteriors under a range of assumptions (Neal, 1992; Roberts & Tweedie, 1996; Durmus & Moulines, 2019). We shall consider the privacy of both the final draw $x^{\mathcal{D}}_{n}\in\mathbb{R}^{d}$, and of the entire chain $(x^{\mathcal{D}}_{1},...,x^{\mathcal{D}}_{n})\in\mathbb{R}^{dn}$, under a non-convex assumption on $U_{\mathcal{D}}$. The particular assumption we place upon the posterior in the following theorem is close to the convexity outside of a ball condition considered in Durmus & Moulines (2017) and Erdogdu et al. (2022). color=red!20]AD: clash of notation for $\mu$ and $c$ to be replaced by $c_{V}$ and $\mathtt{m}$
@@
for all $x,y\in\mathbb{R}^{d}$, and also strongly convex, that is there exists $\mu>0$ such that for all $x,y\in\mathbb{R}^{d}$.
$\langle\nabla K(x)-\nabla K(y),x-y\rangle\geqslant\mu\|x-y\|^{2}.$ (26)
@@
Suppose Assumption 5.1 holds and let $\mathcal{A}_{s}({\mathcal{D}})=x_{n}^{\mathcal{D}}$ be the $n$-th state of the ULA targeting $\pi_{\mathcal{D}}$ with step-size $\gamma\in(0,2\mu/L^{2})$. Then, for $\delta>0$ and $\alpha\geqslant 1$, $\mathcal{A}_{s}({\mathcal{D}})$ is $(\varepsilon_{\delta},\delta)$-differentially private and $(\alpha,\varepsilon_{\alpha})$-Rényi differentially private for
@@
where $C_{3}=c^{2}(\frac{2(L+1)}{\mu-\gamma L^{2}/2}+1)^{2}$.
@@
Note that the above bound does not depend on the number of steps. Furthermore, since recent analysis like (Chewi et al., 2022) suggests that ULA recovers the true posterior in Rényi divergence as $\gamma\to 0$ and $n\to\infty$, taking the limit as $\gamma\to 0$ in Theorem 5.2 suggests that true posterior is differentially private with $C_{3}>0$ replaced with $C_{4}=c(\frac{2(L+1)}{\mu}+1)$.
Now we consider the privacy of the entire chain under slightly different assumptions. color=red!20]AD: $c$ should be $c_{U}$
@@
Suppose Assumption 5.3 holds and let $\mathcal{A}_{p}({\mathcal{D}})=(x^{\mathcal{D}}_{1},\dots,x^{\mathcal{D}}_{n})$ be the path of ULA up to state $n\in\mathbb{N}$. Then, for $\delta>0,\alpha\geqslant 1$, the algorithm $\mathcal{A}_{p}({\mathcal{D}})$ is $(\varepsilon_{\delta},\delta)$-differentially private and $(\alpha,\varepsilon_{\alpha})$-Rényi differentially private for
@@
We use the continuous time interpolation of (25), along with Proposition 4.5. Then since $(x^{\mathcal{D}}_{1},...,x^{\mathcal{D}}_{n})$ is equal in law to a mapping from $(X^{{\mathcal{D}}}_{T})_{t\in[0,n\gamma]}$, the result follows from the data processing inequality (see Theorem 9 in van Erven & Harremos (2014)). ∎
@@ 5.2 Noisy stochastic gradient descent
where $\ell:\mathbb{R}^{d}\times\mathcal{X}\to\mathbb{R}$, $(A_{n})_{n\geqslant 1}$ is a sequence of independent random variables uniformly distributed on subsets of $\{1,...,m\}$ of size $s\leqslant m$, the step size is $\gamma>0$, and $(z_{n})_{n\geqslant 1}$ is a sequence of i.i.d. standard Gaussians on $\mathbb{R}^{d}$, independent of $(A_{n})_{n\geqslant 1}$. Here $\ell(x,d_{i})$ is interpreted as the loss incurred for datum $d_{i}$. We remark that the algorithm (31) is known in the MCMC literature as stochastic gradient Langevin dynamics (Welling & Teh, 2011).
@@
for functions $v:\mathbb{R}^{d}\times\mathcal{S}\to\mathbb{R}$ and $k:\mathbb{R}^{d}\to\mathbb{R}$ that are once continuously differentiable. Furthermore, $\nabla k$ is $L$-Lipschitz and for $c,\mu>0$ one has $\lvert\nabla_{x}v(x,d)\rvert\leqslant c$ for any $x\in\mathbb{R}^{d}$ and any datum $d$, as well as
$\langle\nabla k(x)-\nabla k(y),x-y\rangle\geqslant\mu\lVert x-y\rVert^{2}\;.$ (33)
@@
Suppose Assumption 5.5 holds and consider the algorithm $\mathcal{A}_{s}({\mathcal{D}})=x_{n}^{\mathcal{D}}$ described in (31) with $\gamma\in(0,2\mu/L^{2})$ and stochastic gradient of size $s\leqslant m$. Then, for $\delta>0$ and $\alpha\geqslant 1$, one has that $\mathcal{A}_{s}$ is $(\varepsilon_{\delta},\delta)$-differentially private and $(\alpha,\varepsilon_{\alpha})$-Rényi differentially private for
@@
where $C_{6}=c^{2}\beta(\frac{2(L+1)}{\mu-\gamma L^{2}/2}+1)^{2}$.
@@
The proof here is similar to the proof of Theorem 5.2, with minor alterations due to the stochastic gradient and the inverse temperature parameter $\beta>0$. Full details are given in Appendix C.2. ∎
Theorem 5.6 is only a minor refinement of Theorem 5.2, and in particular the privacy guarantee does not improve with the size $m$ of the dataset. However, in the strongly convex case where each $\nabla v$ is constant, the following theorem (similar to Theorem 2 in Chourasia et al. (2021)) shows that privacy does increase with $s\leqslant m$.
@@
where $C_{7}=\frac{c^{2}\beta}{s^{2}}(\frac{2(L+1)}{\mu-\gamma L^{2}/2}+1)^{2}$.
@@
It is an open problem as to how $m$ and $s$ affect the privacy of the noisy SGD algorithm in the setting of Theorem 5.6. However, the following result on the path of noisy SGD improves with the size $s\leqslant m$ of the stochastic gradient.
@@
The proof is very similar to the proof of Theorem 5.4. Full details are given in Section C.4. ∎
@@ 5.3 Commentary on Bounds Presented
We have studied two kinds of dimension-independent privacy guarantees: for the entire trajectory of an MCMC chain, and for the final draw. In the former case, we improved on known composition bounds. In particular, our $(\varepsilon,\delta)$-DP bounds in Theorems 5.2 and 5.6 feature uniform values of $\delta>0$, and $\varepsilon=O(n+\sqrt{n\log(1/\delta)})$ dependence on the number of steps $n$, whilst the advanced composition bounds presented in Kairouz et al. (2015) achieve $\varepsilon=O(n+\sqrt{n\log(e+(\varepsilon\sqrt{n}/\delta))})$. This essentially matches bounds presented in Chourasia et al. (2021). On the other hand, our bounds for the privacy of the final draw in Theorems 5.2 and 5.6 are uniform in time in a non-convex setting on an unbounded space, which answers the open problem of question Question 1.1 in Altschuler & Talwar (2022). Also, unlike the results of Altschuler & Talwar (2022), our bounds do not depend on the size of the state space (which is unbounded), and do not blow up as the step-size goes to 00. In this sense, we therefore generalise the results of Chourasia et al. (2021) to a non-convex setting. Indeed, our results match Chourasia et al. (2021) in the strongly convex regime.
@@ Acknowledgements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proofs and general results for Section 3
In this section we give proofs and general statements relative to Section 3. These rely on the following definition.
@@ Appendix A Proofs and general results for Section 3
A.1 Auxiliary result for Proposition 3.2
@@
Let $\{\mu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ and $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ that satisfy $\lVert\mu_{\mathcal{D}}-\nu_{\mathcal{D}}\rVert_{\mathrm{TV}}\leqslant\beta$ for any ${\mathcal{D}}\in\mathcal{S}$. If $\mu$ is $(\varepsilon,\delta)$-differentially private, then $\nu$ is $(\varepsilon,\delta+\beta(e^{\varepsilon}+1))$-differentially private.
@@
A.2 General result for Proposition 3.4
@@
Let $\{\mu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ and $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ be two families of probability distributions. Assume $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ is $(\varepsilon,\delta_{\nu})$-differentially private and $\{\mu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ is not $(\varepsilon,\delta_{\mu})$-DP for $\delta_{\mu}>\delta_{\nu}$, i.e. there exists at least one pair of adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ for which there is a measurable set $\mathsf{B}$ such that
@@
[…]
@@
We take any two adjacent datasets $({\mathcal{D}},{\mathcal{D}}^{\prime})\in\overline{\mathcal{S}}$. Notice that (44) implies that
@@
$\mu_{\mathcal{D}}(\mathsf{B})-\nu_{\mathcal{D}}(\mathsf{B})$ $\leqslant-(e^{-\varepsilon}\delta_{\mu}-e^{-\varepsilon}\delta_{\nu})+e^{-\varepsilon}\mu_{{\mathcal{D}}^{\prime}}(\mathsf{B})-e^{-\varepsilon}\nu_{{\mathcal{D}}^{\prime}}(\mathsf{B})$ (49)
$\leqslant-(e^{-\varepsilon}\delta_{\mu}-e^{-\varepsilon}\delta_{\nu})+e^{-\varepsilon}(\mu_{{\mathcal{D}}^{\prime}}(\mathsf{B})-\nu_{{\mathcal{D}}^{\prime}}(\mathsf{B}))\;.$ (50)
Suppose now that $\lVert\mu_{{\mathcal{D}}^{\prime}}-\nu_{{\mathcal{D}}^{\prime}}\rVert_{\mathrm{TV}}\leqslant\zeta$ for some constant $0<\zeta<\delta_{\mu}-\delta_{\nu}$. This implies that $\mu_{{\mathcal{D}}^{\prime}}(\mathsf{B})-\nu_{{\mathcal{D}}^{\prime}}(\mathsf{B})\leqslant\zeta$. Therefore, we find
@@
Since the right hand side is negative by construction, taking the absolute value we find
$\lvert\mu_{\mathcal{D}}(\mathsf{B})-\nu_{\mathcal{D}}(\mathsf{B})\rvert\geqslant e^{-\varepsilon}(\delta_{\mu}-\delta_{\nu}-\zeta)\;.$ (52)
By the definition of TV distance we find
@@
This shows that there always exists a dataset $\tilde{\mathcal{D}}\in\{{\mathcal{D}},{\mathcal{D}}^{\prime}\}$ such that
@@
Optimising the bound for $\zeta\in(0,\delta_{\mu}-\delta_{\nu})$ we find $\zeta^{*}=\frac{e^{-\varepsilon}}{1+e^{-\varepsilon}}(\delta_{\mu}-\delta_{\nu}),$ which satisfies the equation $\zeta^{*}=e^{-\varepsilon}(\delta_{\mu}-\delta_{\nu}-\zeta^{*}).$
A.3 General result for Proposition 3.7
@@
where $g_{\mathcal{D}}$ is a random variable that depends on the dataset, and $L$ satisfies Assumption 3.5 and is independent of $g_{\mathcal{D}}$. We make the following assumptions, which is the general version of Assumption 3.6.
@@
Let ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ denote two adjacent datasets. There exist $\eta>0,\delta<1$ that are independent of ${\mathcal{D}},{\mathcal{D}}^{\prime}$ such that
@@
Assumption 3.5 and Assumption A.4 are enough to obtain the following guarantee on the differential privacy of the randomised algorithm (55).
@@
Let $\mathcal{A}$ be the randomised mechanism defined in (55), where $L$ is independent of $g_{\mathcal{D}}$ for any ${\mathcal{D}}\in\mathcal{S}.$ Suppose Assumption A.4 holds, as well as Assumption 3.5 for $\eta$ as in Assumption A.4. Then $\mathcal{A}$ is $(\varepsilon,\tilde{\delta}+\delta)$-differentially private.
@@
The first term on the right hand side of (57) can be bounded using Assumption 3.5. Indeed, Assumption 3.5 can be used since we are on the event $\left\lvert g_{\mathcal{D}}-g_{{\mathcal{D}}^{\prime}}\right\rvert\leqslant\eta$ and because $L$ and $g_{{\mathcal{D}}},g_{{\mathcal{D}}^{\prime}}$ are independent. Following this reasoning, we find
@@
In the last inequality we simply discarded the event $\{\left\lvert g_{\mathcal{D}}-g_{{\mathcal{D}}^{\prime}}\right\rvert\leqslant\eta\}.$ Applying Assumption A.4 we can bound the second term in (57) as follows:

### v2

@@ 5 Privacy of Langevin-based algorithms in the non-convex setting
In this section we obtain (Rényi)-DP for both the trajectory and the final value of two Langevin-based MCMC algorithms: ULA and noisy SGD. All bounds in this section are derived from Propositions˜4.6 and 4.5.
@@ 5.1 Sampling from Bayesian posteriors with ULA
with initial condition $x^{\mathcal{D}}_{0}=x_{0}\in\mathbb{R}^{d},$ and where $(z_{n})_{n\geqslant 1}$ is a sequence of i.i.d. standard Gaussians on $\mathbb{R}^{d}$, and $\gamma>0$ is the step size. ULA arises as the Euler discretisation of the overdamped Langevin diffusion and has been shown to be successful for sampling from Bayesian posteriors under a range of assumptions (Neal, 1992; Roberts & Tweedie, 1996; Durmus & Moulines, 2019). We shall consider the privacy of both the final draw $x^{\mathcal{D}}_{n}\in\mathbb{R}^{d}$, and of the entire chain $(x^{\mathcal{D}}_{1},...,x^{\mathcal{D}}_{n})\in\mathbb{R}^{dn}$, under a non-convex assumption on $U_{\mathcal{D}}$. The particular assumption we place upon the posterior in the following theorem is close to the convexity outside of a ball condition considered in Durmus & Moulines (2017) and Erdogdu et al. (2022). The strongly convex part $K$ could be interpreted as a regulariser.
@@
for all $x,y\in\mathbb{R}^{d}$, and also strongly convex, that is there exists $a>0$ such that for all $x,y\in\mathbb{R}^{d}$.
$\langle\nabla K(x)-\nabla K(y),x-y\rangle\geqslant a\|x-y\|^{2}.$ (28)
@@
Suppose Assumption˜5.1 holds and let $\mathcal{A}_{s}({\mathcal{D}})=x_{n}^{\mathcal{D}}$ be the $n$-th state of the ULA targeting $\pi_{\mathcal{D}}$ with step-size $\gamma\in(0,2a/L^{2})$. Then, for $\delta>0$ and $\alpha\geqslant 1$, $\mathcal{A}_{s}({\mathcal{D}})$ is $(\varepsilon_{\delta},\delta)$-differentially private and $(\alpha,\varepsilon_{\alpha})$-Rényi differentially private for
@@
where $C_{3}=c^{2}(\frac{2(L+1)}{a-\gamma L^{2}/2}+1)^{2}$.
@@
Note that the above bound does not depend on the number of steps. Furthermore, since recent analysis like (Chewi et al., 2022) suggests that ULA recovers the true posterior in Rényi divergence as $\gamma\to 0$ and $n\to\infty$, taking the limit as $\gamma\to 0$ in Theorem 5.2 suggests that true posterior is differentially private with $C_{3}>0$ replaced with $C_{4}=c(\frac{2(L+1)}{a}+1)$.
Now we consider the privacy of the entire chain under slightly different assumptions.
@@
Suppose Assumption˜5.3 holds and let $\mathcal{A}_{p}({\mathcal{D}})=(x^{\mathcal{D}}_{1},\dots,x^{\mathcal{D}}_{n})$ be the path of ULA up to state $n\in\mathbb{N}$. Then, for $\delta>0,\alpha\geqslant 1$, the algorithm $\mathcal{A}_{p}({\mathcal{D}})$ is $(\varepsilon_{\delta},\delta)$-differentially private and $(\alpha,\varepsilon_{\alpha})$-Rényi differentially private for
@@
We use the continuous time interpolation of (27), along with Proposition˜4.6. Then since $(x^{\mathcal{D}}_{1},...,x^{\mathcal{D}}_{n})$ is equal in law to a mapping from $(X^{{\mathcal{D}}}_{T})_{t\in[0,n\gamma]}$, the result follows from the data processing inequality (see Theorem 9 in van Erven & Harremos (2014)). ∎
@@ 5.2 Noisy stochastic gradient descent
where $\ell:\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}$, $(A_{n})_{n\geqslant 1}$ is a sequence of independent random variables uniformly distributed on subsets of $\{1,...,m\}$ of size $s\leqslant m$, the step size is $\gamma>0$, and $(z_{n})_{n\geqslant 1}$ is a sequence of i.i.d. standard Gaussians on $\mathbb{R}^{d}$, independent of $(A_{n})_{n\geqslant 1}$. Here $\ell(x,d_{i})$ is interpreted as the loss incurred for datum $d_{i}$. We remark that the algorithm (33) is known in the MCMC literature as stochastic gradient Langevin dynamics (Welling & Teh, 2011).
@@
for functions $v:\mathbb{R}^{d}\times\mathcal{S}\to\mathbb{R}$ and $k:\mathbb{R}^{d}\to\mathbb{R}$ that are once continuously differentiable. Furthermore, $\nabla k$ is $L$-Lipschitz and for $c,a>0$ one has $\lvert\nabla_{x}v(x,d)\rvert\leqslant c$ for any $x\in\mathbb{R}^{d}$ and any datum $d$, as well as
$\langle\nabla k(x)-\nabla k(y),x-y\rangle\geqslant a\lVert x-y\rVert^{2}\;.$ (35)
@@
Suppose Assumption˜5.5 holds and consider the algorithm $\mathcal{A}_{s}({\mathcal{D}})=x_{n}^{\mathcal{D}}$ described in (33) with $\gamma\in(0,2a/L^{2})$ and stochastic gradient of size $s\leqslant m$. Then, for $\delta>0$ and $\alpha\geqslant 1$, one has that $\mathcal{A}_{s}$ is $(\varepsilon_{\delta},\delta)$-differentially private and $(\alpha,\varepsilon_{\alpha})$-Rényi differentially private for
@@
where $C_{6}=c^{2}\beta(\frac{2(L+1)}{a-\gamma L^{2}/2}+1)^{2}$.
@@
The proof here is similar to the proof of Theorem˜5.2, with minor alterations due to the stochastic gradient and the inverse temperature parameter $\beta>0$. Full details are given in Appendix C.2. ∎
Theorem˜5.6 is only a minor refinement of Theorem 5.2, and in particular the privacy guarantee does not improve with the size $m$ of the dataset. However, in the strongly convex case where each $\nabla v$ is constant, the following theorem (similar to Theorem 2 in Chourasia et al. (2021)) shows that privacy does increase with $s\leqslant m$.
@@
where $C_{7}=\frac{c^{2}\beta}{s^{2}}(\frac{2(L+1)}{a-\gamma L^{2}/2}+1)^{2}$.
@@
It is an open problem as to how $m$ and $s$ affect the privacy of the noisy SGD algorithm in the setting of Theorem˜5.6. However, the following result on the path of noisy SGD improves with the size $s\leqslant m$ of the stochastic gradient.
@@
The proof is very similar to the proof of Theorem˜5.4. Full details are given in Section˜C.4. ∎
@@ 5.3 Commentary on Bounds Presented
We have studied two kinds of dimension-independent privacy guarantees: for the final draw, and for the entire trajectory of an MCMC chain. In particular, our bounds for the privacy of the final draw in Theorems˜5.2 and 5.6 are uniform-in-time in a non-convex setting on an unbounded space, which addresses Question 1.2 and provides an upper bound for Question 1.1 in Altschuler & Talwar (2022). Also, unlike the results of Altschuler & Talwar (2022), our bounds do not depend on the size of the state space (which is unbounded), and do not blow up as the step-size goes to 0. In this sense, we therefore generalise the results of Chourasia et al. (2021) to a non-convex setting. Indeed, our results match Chourasia et al. (2021) in the strongly convex regime.
On the other hand, our results for the entire trajectory we improved on known composition bounds for $(\epsilon,\delta)$-privacy, matching composition bounds for Rényi privacy. This removes the need for complicated $(\epsilon,\delta)$-DP privacy accounting, and provides a simple framework which extends naturally to the continuous time case. In particular, our $(\varepsilon,\delta)$-DP bounds in Theorems˜5.2 and 5.6 feature uniform values of $\delta>0$, and $\varepsilon=O(n+\sqrt{n\log(1/\delta)})$ dependence on the number of steps $n$, whilst the advanced composition bounds presented in Kairouz et al. (2015) achieve $\varepsilon=O(n+\sqrt{n\log(e+(\varepsilon\sqrt{n}/\delta))})$.
@@ Acknowledgements
This paper presents work whose goal is to advance the field of Machine Learning. We believe by extending the theoretical study of differential privacy, practitioners will be able to design methods that better protect the privacy of individuals.
Appendix A Proofs and general results for Section˜3
In this section we give proofs and general statements relative to Section˜3. These rely on the following definition.
@@ Appendix A Proofs and general results for Section˜3
A.1 Auxiliary result for Proposition˜3.2
@@
Let $\{\mu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ and $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ be two families of probability distributions that satisfy $\lVert\mu_{\mathcal{D}}-\nu_{\mathcal{D}}\rVert_{\mathrm{TV}}\leqslant\beta$ for any ${\mathcal{D}}\in\mathcal{S}$. If $\{\mu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ is $(\varepsilon,\delta)$-differentially private, then $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ is $(\varepsilon,\delta+\beta(e^{\varepsilon}+1))$-differentially private.
@@
A.2 General result for Proposition˜3.3
@@
Let $\{\mu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ and $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ be two families of probability distributions that satisfy $\lVert\mu_{\mathcal{D}}-\nu_{\mathcal{D}}\rVert_{\mathrm{TV}}\leqslant\beta$ for any ${\mathcal{D}}\in\mathcal{S}$. Let $\varepsilon>0$, $\delta\in[(1+e^{\varepsilon})\beta,1)$, and suppose $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ is not $(\varepsilon,\delta)$-differentially private. Then, $\{\mu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ is not $(\varepsilon,\delta-(1+e^{\varepsilon})\beta)$-differentially private.
Proof.
Let ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ denote two adjacent datasets and $\mathsf{B}$ be a measurable set such that
$\nu_{{\mathcal{D}}^{\prime}}(\mathsf{B})\geqslant e^{\varepsilon}\nu_{\mathcal{D}}(\mathsf{B})+\delta,$ (46)
i.e. such that the definition of DP for the family $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ is violated. Now, applying the inequality (42) followed by the inequality above we find
$\mu_{{\mathcal{D}}^{\prime}}(\mathsf{B})$ $\geqslant\nu_{{\mathcal{D}}^{\prime}}(\mathsf{B})-\beta$ (47)
$\geqslant e^{\varepsilon}\nu_{\mathcal{D}}(\mathsf{B})+\delta-\beta.$ (48)
Then, we apply inequality (41) to find
$\mu_{{\mathcal{D}}^{\prime}}(\mathsf{B})$ $\geqslant e^{\varepsilon}\mu_{\mathcal{D}}(\mathsf{B})+\delta-(1+e^{\varepsilon})\beta,$ (49)
which proves the result. ∎
A.3 General result for Proposition˜3.4
Proposition A.4.
Let $\{\mu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ and $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ be two families of probability distributions. Assume $\{\nu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ is $(\varepsilon,\delta_{\nu})$-differentially private and $\{\mu_{\mathcal{D}}:{\mathcal{D}}\in\mathcal{S}\}$ is not $(\varepsilon,\delta_{\mu})$-DP for $\delta_{\mu}>\delta_{\nu}$, i.e. there exists at least one pair of adjacent datasets ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ and a measurable set $\mathsf{B}$ such that
@@
Then, for any $\{{\mathcal{D}},{\mathcal{D}}^{\prime}\}\in\overline{\mathcal{S}}$ there exists $\tilde{\mathcal{D}}\in\{{\mathcal{D}},{\mathcal{D}}^{\prime}\}$ such that
@@
We take any pair of datasets $\{{\mathcal{D}},{\mathcal{D}}^{\prime}\}\in\overline{\mathcal{S}}$. Notice that (50) implies that
@@
$\mu_{\mathcal{D}}(\mathsf{B})-\nu_{\mathcal{D}}(\mathsf{B})$ $<-(e^{-\varepsilon}\delta_{\mu}-e^{-\varepsilon}\delta_{\nu})+e^{-\varepsilon}\mu_{{\mathcal{D}}^{\prime}}(\mathsf{B})-e^{-\varepsilon}\nu_{{\mathcal{D}}^{\prime}}(\mathsf{B})$ (55)
$<-e^{-\varepsilon}(\delta_{\mu}-\delta_{\nu})+e^{-\varepsilon}(\mu_{{\mathcal{D}}^{\prime}}(\mathsf{B})-\nu_{{\mathcal{D}}^{\prime}}(\mathsf{B}))\;.$ (56)
Suppose now that $\lVert\mu_{{\mathcal{D}}^{\prime}}-\nu_{{\mathcal{D}}^{\prime}}\rVert_{\mathrm{TV}}\leqslant\zeta$ for some constant $0\leqslant\zeta\leqslant\delta_{\mu}-\delta_{\nu}$. This implies that $\mu_{{\mathcal{D}}^{\prime}}(\mathsf{B})-\nu_{{\mathcal{D}}^{\prime}}(\mathsf{B})\leqslant\zeta$. Therefore, we find
@@
Since $\delta_{\mu}-\delta_{\nu}-\zeta\geqslant 0$ by construction, taking the absolute value we find
$\lvert\mu_{\mathcal{D}}(\mathsf{B})-\nu_{\mathcal{D}}(\mathsf{B})\rvert>e^{-\varepsilon}(\delta_{\mu}-\delta_{\nu}-\zeta)\;.$ (58)
By the definition of TV distance
@@
Clearly, when $\lVert\mu_{{\mathcal{D}}^{\prime}}-\nu_{{\mathcal{D}}^{\prime}}\rVert_{\mathrm{TV}}\leqslant\zeta$ does not hold, it must be that $\lVert\mu_{{\mathcal{D}}^{\prime}}-\nu_{{\mathcal{D}}^{\prime}}\rVert_{\mathrm{TV}}>\zeta$. This shows that there exists a dataset $\tilde{\mathcal{D}}\in\{{\mathcal{D}},{\mathcal{D}}^{\prime}\}$ such that either $\lVert\mu_{\tilde{\mathcal{D}}}-\nu_{\tilde{\mathcal{D}}}\rVert_{\mathrm{TV}}>\zeta$ or $\lVert\mu_{\tilde{\mathcal{D}}}-\nu_{\tilde{\mathcal{D}}}\rVert_{\mathrm{TV}}>e^{-\varepsilon}(\delta_{\mu}-\delta_{\nu}-\zeta)$. Therefore it always holds that
@@
Optimising the bound for $\zeta\in[0,\delta_{\mu}-\delta_{\nu}]$, we find that the lower bound above is maximised for $\zeta^{*}=\frac{e^{-\varepsilon}}{1+e^{-\varepsilon}}(\delta_{\mu}-\delta_{\nu}),$ which satisfies the equation $\zeta^{*}=e^{-\varepsilon}(\delta_{\mu}-\delta_{\nu}-\zeta^{*}).$
A.4 General result for Proposition˜3.7
@@
where $g_{\mathcal{D}}$ is a random variable that depends on the dataset, and $L$ satisfies Assumption˜3.5 and is independent of $g_{\mathcal{D}}$. We make the following assumptions, which is the general version of Assumption˜3.6.
@@
Let ${\mathcal{D}},{\mathcal{D}}^{\prime}\in\mathcal{S}$ denote two adjacent datasets. There exist $\eta>0,\tilde{\delta}<1$ that are independent of ${\mathcal{D}},{\mathcal{D}}^{\prime}$ such that
@@
Assumption˜3.5 and Assumption˜A.5 are enough to obtain the following guarantee on the differential privacy of the randomised algorithm (61).
@@
Let $\mathcal{A}$ be the randomised mechanism defined in (61), where $L$ is independent of $g_{\mathcal{D}}$ for any ${\mathcal{D}}\in\mathcal{S}.$ Suppose Assumption˜A.5 holds, as well as Assumption˜3.5 for $\eta$ as in Assumption˜A.5. Then $\mathcal{A}$ is $(\varepsilon,\tilde{\delta}+\delta)$-differentially private.
@@
The first term on the right hand side of (63) can be bounded using Assumption˜3.5. Indeed, Assumption˜3.5 can be used since we are on the event $\left\lvert g_{\mathcal{D}}-g_{{\mathcal{D}}^{\prime}}\right\rvert\leqslant\eta$ and because $L$ and $g_{{\mathcal{D}}},g_{{\mathcal{D}}^{\prime}}$ are independent. Following this reasoning, we find
@@
In the last inequality we simply discarded the event $\{\left\lvert g_{\mathcal{D}}-g_{{\mathcal{D}}^{\prime}}\right\rvert\leqslant\eta\}.$ Applying Assumption˜A.5 we can bound the second term in (63) as follows: